# Geometric Time-Dependent Density Functional Theory

### v1

@@ Abstract
For systems close to equilibrium, standard adiabatic TDDFT approximations work well. However, this is not the case far from equilibrium. It is thus natural to use our new geometric theory to correct the limitations of the existing adiabatic approximations. For any given approximation $V^{\rm app}[\Psi_{0},\Phi_{0},\rho]$ of the exact standard Kohn–Sham Hxc potential, we thus define the corresponding Kohn–Sham equation with geometric correction as
$i\partial_{t}\varphi_{n}(t,\mathbf{r})=\Big(-\frac{\nabla^{2}}{2}+V_{\rm ext}(t,\mathbf{r})+V^{\rm app}[\Psi_{0},\Phi_{0},\rho_{\Phi}](t,\mathbf{r})\\ +i\big[W^{\rm corr}(t)\,,\,\gamma_{\Phi(t)}\big]\Big)\varphi_{n}(t,\mathbf{r}).$ (17)
@@
We can repeat everything we said before in the absence of $V^{\rm app}$. For instance, this equation will reproduce the exact Schrödinger density $\rho^{\rm S}$ if $W^{\rm corr}=\widetilde{W}[\Psi_{0},\Phi]$, the functional defined in (14), because the continuity equation is insensitive to external local potentials. We can similarly use functionals depending only on $\rho$ such as $W^{\rm GKS}$ in (15) or $\widetilde{W}^{\rm GKS}$ in (16), but the definitions of $W_{0}$, $\mathbf{j}_{0}^{\rm GP}$ and $\Phi^{\rm GP}$ have to be modified to take into account the new nonlinear potential $V^{\rm app}$. If $V^{\rm app}$ is a good adiabatic approximation, we expect $W^{\rm corr}$ to be very small when close to a ground state and to only be significant in non-adiabatic situations. In the next paragraph we consider the case of two-electron spin-singlet states, for which Eq. (17) can be solved explicitly if $V_{\rm ext}+V^{\rm app}$ is the exact adiabatic Kohn–Sham potential.
@@
In standard time-dependent Kohn–Sham theory, $W$ is taken to vanish, $\theta$ is obtained by solving (18a), and $V$ is given by (18b) with that $\theta$. For the geometric Kohn–Sham equation (13) where $V=V_{\rm ext}$ the situation is reversed: One should first solve (18b) to find $\theta$ and then $W$ is given by (18a). There are many other solutions to (18). An important observation is that, all the solutions can be parametrized by $\theta$ leading to explicit formulas for $V$ and $W$ from (18). The simplest case is $\theta\equiv 0$ with
@@
The potential $V^{\rm eaKS}$ is the exact adiabatic Kohn–Sham potential, defined as the potential for which the time-dependent Kohn–Sham system has ground-state density $\rho^{\rm S}$. The term $W^{\text{na}}$ is the corresponding non-adiabatic geometric correction appearing in (17) for that case. It is remarkable that $W^{\text{na}}$ describing all the non-adiabatic effects is such a simple and explicit functional of $\rho^{\rm S}$.
@@
We next consider a two-electron singlet in 1D, with a soft-Coulomb interaction $V_{\text{ee}}(x)=1/\sqrt{x^{2}+1}$. This framework was extensively studied [34, 35, 36, 28, 37, 38], because standard adiabatic TDDFT approximations fail to reproduce, even qualitatively, the exact time-dependent density for this 1D toy model in much the same way as they do for real 3D Coulombic systems. We consider two different time-dependent external potentials $V_{\rm ext}^{(1)}$ and $V_{\rm ext}^{(2)}$.
First, as in [34, 38], we study the Rabi oscillations in a Helium-like atom perturbed by an electric dipole interaction, corresponding to
@@
To estimate $\rho^{\rm S}$, we follow common practice and solve a two-level time-dependent Schrödinger equation in the basis $(\Psi_{\rm g},\Psi_{\rm e})$. These two functions are obtained by diagonalizing a finite-difference approximation of the two-electron Hamiltonian at time $t=0$. The non-adiabatic contribution $V^{\text{na}}$ to the time-dependent exact Kohn–Sham potential is computed from (18) with $W=0$. In 1D, we have the explicit formula [39, 40]
@@
Figure 2: Rabi oscillations in a 1D Helium-like atom with soft-Coulomb potential subjected to a time-dependent uniform electric field with amplitude $E=0.00667$ and angular frequency $\omega=0.5336$. Heatmaps of the non-adiabatic contribution $V^{\text{na}}$ to the exact time-dependent Kohn-Sham potential (left) and of the non-adiabatic geometric correction term $W^{\text{na}}$ (right), at times $|t-T_{\rm R}/4|\leqslant 2T_{\rm opt}$ (bottom) and $|t-T_{\rm R}/2|\leqslant 2T_{\rm opt}$ (top) with $T_{\rm opt}=2\pi/\omega$ the optical period.
Figure 3: Charge transfer in 1D between a Helium-like atom and a model of a closed-shell atom with frozen electrons with soft-Coulomb potential obtained by a time-dependent uniform electric field with amplitude $E=0.00667$ and angular frequency $\omega=0.112$. Heatmaps of the non-adiabatic contribution $V^{\text{na}}$ to the exact time-dependent Kohn-Sham potential (left) and of the non-adiabatic geometric correction term $W^{\text{na}}$ (right), at times $|t-T_{\rm R}/8|\leqslant 2T_{\rm opt}$ (bottom) and $|t-T_{\rm R}/4|\leqslant 2T_{\rm opt}$ (top) with $T_{\rm opt}=2\pi/\omega$ the optical period.
@@ Conclusion.
We have proposed a new formulation of TDDFT based on geometric considerations. Our numerical simulations in the two-electron singlet case suggest that approximating the non-adiabatic correction $W^{\rm na}$ could be easier than the corresponding non-adiabatic part of the standard Kohn–Sham potential, because it is a much smoother function of both time and space. Our new formulation could thus be more adapted to the description of out-of-equilibrium processes. In the companion article [10] we study the mathematical foundations of our geometric approach for finite-dimensional systems. Future work will be devoted to constructing practical approximations of the geometric correction $W$ for real systems.
@@ Acknowledgments.
This work has benefited from French State support managed by ANR under the France 2030 program through the MaQuI CNRS Risky and High-Impact Research programme (RI)2 (grant agreement ANR-24-RRII-0001).
@@ Appendix A Geometric Runge–Gross theorem
The Runge–Gross theorem [9] asserts that two time-dependent external potentials giving rise to the same density must necessarily be equal, up to a time-dependent additive constant. This is a fundamental result in TDDFT. Thus far, the mathematical proof of this result has only been possible under the assumption that the Schrödinger wavefunction and the time-dependent potentials are analytic in time [42]. The Runge–Gross argument therefore only applies to very smooth potentials such as the soft-Coulomb potential. Here we provide a rigorous proof that this also works for our geometric theory, following the lines of [42]. More precisely, we work in space dimension $d\in\{1,2,3\}$, and given a time-dependent Hamiltonian $\widehat{H}(t)$ of the form (1) and two functions $W_{1}(t,x)$ and $W_{2}(t,x)$, we assume that
  1. (i)
@@
  2. (ii)
    the external potential $V_{\rm ext}(t,\mathbf{r})$ and the functions $W_{1}(t,\mathbf{r})$ and $W_{2}(t,\mathbf{r})$ are infinitely differentiable in space-time, with bounded space derivatives for all $t$;
  3. (iii)
    for any $\mathbf{r}\in\mathbb{R}^{d}$, the function $t\mapsto W_{2}(t,\mathbf{r})-W_{1}(t,\mathbf{r})$ is real-analytic;
  4. (iv)
    the initial state $\Psi_{0}$ is infinitely differentiable in space and all its derivatives are square-integrable;
  5. (v)
    the initial state $\Psi_{0}$ satisfies the unique $V$-representability property
@@
The equalities in (20) are understood Lebesgue-almost everywhere. The unique $V$-representability property (20) is further studied in [43, 10]. This is the main property used in the proof of the Hohenberg–Kohn theorem for ground-state DFT, where it follows from the unique continuation principle [44, 45]. In particular, (20) is valid if $\Psi_{0}$ is the ground state of a Hamiltonian of the above form. For the theorem we do not need that $\Psi_{0}$ is a ground state but we do assume the condition (20). As explained in [10], this is the main property necessary for the geometric interpretation of TDDFT.
@@
$i\partial_{t}\Psi_{m}(t)=\bigg(\widehat{H}(t)+i\sum_{j=1}^{N}W_{m}(t,\mathbf{r}_{j})\bigg)\Psi_{m}(t)$
with $m=1$ or $m=2$, starting at the same initial condition $\Psi_{0}$. If $\rho_{\Psi_{1}(t)}=\rho_{\Psi_{2}(t)}$ for $t\in[0,T)$, then $W_{1}(t,\mathbf{r})=W_{2}(t,\mathbf{r})$ for all $\mathbf{r}\in\mathbb{R}^{d}$ and all $t\in[0,T)$.
The proof of Theorem 1 follows [42]. We show that
@@
for all $\ell$ and all $\mathbf{r}\in\mathbb{R}^{d}$. The conclusion then follows from the assumed analyticity in time. The smoothness of the potentials and $\Psi_{0}$ implies that we can freely differentiate Schrödinger’s equation in time [42]. We denote by $\widehat{W}_{m}(t):=\sum_{j=1}^{N}W_{m}(t,\mathbf{r}_{j})$ the many-body potentials associated with $W_{1}$ and $W_{2}$. We also introduce the relative potential $W:=W_{2}-W_{1}$ which we want to prove vanishes, and the corresponding $\widehat{W}:=\widehat{W}_{2}-\widehat{W}_{1}$. We call $\rho(t,\mathbf{r})$ the common density of $\Psi_{1}(t)$ and $\Psi_{2}(t)$.
We give ourselves a smooth time-independent potential $V(\mathbf{r})$, that we will specify later. Using again the notation $\widehat{V}:=\sum_{j=1}^{N}V(\mathbf{r}_{j})$, we have
$\int_{\mathbb{R}^{d}}V(\mathbf{r})\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{V}\Psi_{m}(t)\right\rangle}.$
@@
$\int_{\mathbb{R}^{d}}V(\mathbf{r})\partial_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}\\ =\Big\langle\Psi_{m}(t),\big(i[\widehat{T},\widehat{V}]+\widehat{V}\widehat{W}_{m}(t)+\widehat{W}_{m}(t)\widehat{V}\big)\Psi_{m}(t)\Big\rangle.$ (22)
We evaluate (22) at time $t=0$ and recall that $\Psi_{1}(0)=\Psi_{2}(0)=\Psi_{0}$. Subtracting the two equations leads to
${\left\langle\Psi_{0},(\widehat{V}\widehat{W}(0)+\widehat{W}(0)\widehat{V})\Psi_{0}\right\rangle}=0$
where we recall that $W(0,\mathbf{r})=W_{2}(0,\mathbf{r})-W_{1}(0,\mathbf{r})$. It is natural to take $V(\mathbf{r})=W(0,\mathbf{r})$ so that the previous relation reduces to $\|\widehat{W}(0)\Psi_{0}\|^{2}=0$, that is, $\widehat{W}(0)\Psi_{0}=0$. From the unique $V$-representability property (20), this implies $W(0,\mathbf{r})=W_{2}(0,\mathbf{r})-W_{1}(0,\mathbf{r})=0$ almost everywhere, hence everywhere by continuity. This is the claimed result (21) for $\ell=0$.
Next we go on and differentiate once more (22), still with a general potential $V$. To simplify the expression, we denote by $\widehat{L}_{m}^{(1)}(t):=i[\widehat{T},\widehat{V}]+\widehat{V}\widehat{W}_{m}(t)+\widehat{W}_{m}(t)\widehat{V}$ the Hermitian operator appearing in (22). The second-order derivative equals
$\int_{\mathbb{R}^{d}}V(\mathbf{r})\partial^{2}_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{K}^{(1)}_{m}(t)\Psi_{m}(t)\right\rangle}\\ +{\left\langle\Psi_{m}(t),\big(\widehat{V}\partial_{t}\widehat{W}_{m}(t)+\partial_{t}\widehat{W}_{m}(t)\widehat{V}\big)\Psi_{m}(t)\right\rangle}$ (23)
@@
${\left\langle\Psi_{0},\big(\widehat{V}\partial_{t}\widehat{W}(0)+\partial_{t}\widehat{W}(0)\widehat{V}\big)\Psi_{0}\right\rangle}=0.$
Taking $V(\mathbf{r})=\partial_{t}W(0,\mathbf{r})=\partial_{t}W_{2}(0,\mathbf{r})-\partial_{t}W_{1}(0,\mathbf{r})$ and using again (20) for that $V$, we arrive at the conclusion that $\partial_{t}W_{2}(0,\mathbf{r})=\partial_{t}W_{1}(0,\mathbf{r})$, which is (21) for $\ell=1$.
@@
$\int_{\mathbb{R}^{d}}V(\mathbf{r})\partial^{\ell}_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{K}^{(\ell-1)}_{m}(t)\Psi_{m}(t)\right\rangle}\\ +{\left\langle\Psi_{m}(t),\big(\widehat{V}\partial^{\ell}_{t}\widehat{W}_{m}(t)+\partial^{\ell}_{t}\widehat{W}_{m}(t)\widehat{V}\big)\Psi_{m}(t)\right\rangle}$ (24)
@@
$\widehat{L}^{(\ell-1)}_{m}(t):=\widehat{K}^{(\ell-1)}_{m}(t)+\widehat{V}\partial^{\ell-1}_{t}\widehat{W}_{m}(t)+\partial^{\ell-1}_{t}\widehat{W}_{m}(t)\widehat{V}$.

### v2

@@ Abstract
For systems close to equilibrium, standard adiabatic TDDFT approximations work well. However, this is not the case far from equilibrium. It is thus natural to use our new geometric theory to correct the limitations of the existing adiabatic approximations. For any given approximation $V^{\rm app}_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho]$ to the exact Hxc potential in Eq. (16), we define the corresponding Kohn–Sham scheme with geometric correction as
$i\partial_{t}\varphi_{n}(t,\mathbf{r})=\Big(-\frac{\nabla^{2}}{2}+V_{\rm ext}(t,\mathbf{r})+V^{\rm app}_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho_{\Phi}](t,\mathbf{r})\\ +i\big[W^{\rm corr}(t)\,,\,\gamma_{\Phi(t)}\big]\Big)\varphi_{n}(t,\mathbf{r}).$ (19)
@@
We can repeat everything we said before in the absence of $V^{\rm app}_{\rm Hxc}$. Namely, Eq. 19 will reproduce the exact Schrödinger density $\rho^{\rm S}$ if we take $W^{\rm corr}=\widetilde{W}^{\rm KS}$ with $V_{\rm ext}$ replaced by $V_{\rm ext}+V^{\rm app}_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho]$ inside $W_{0}$ in the right-hand side of Eq. (17) or $W^{\rm corr}=W^{\rm KS}$ with $V[\Psi_{0},\rho]$ replaced by $V[\Psi_{0},\rho]+V^{\rm app}_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho]$ in Eq. (18). If $V^{\rm app}_{\rm Hxc}$ is a good adiabatic approximation, we expect $W^{\rm corr}$ to be very small when close to a ground state and to only be significant in non-adiabatic situations. In the next paragraph we consider the case of two-electron spin-singlet states, for which Eq. (19) can be solved explicitly if $V_{\rm ext}+V^{\rm app}_{\rm Hxc}$ is the exact adiabatic Kohn–Sham potential for $\rho^{\rm S}$.
@@
In standard time-dependent Kohn–Sham theory where $W=0$, $\theta$ is obtained by solving (20a), and $V$ is given by (20b) with that $\theta$. For the geometric Kohn–Sham equation (15) where $V=V_{\rm ext}$ the situation is reversed: One should first solve (20b) to find $\theta$ and then $W$ is given by (20a). Note in passing that for the universal functional of Eq. (18) we take $V=V[\Psi_{0},\rho]$ and solve (20) with a general $\rho$ in place of $\rho^{\rm S}$. There are many solutions to (20). An important observation is that they can be parametrized by $\theta$, leading to explicit formulas for $V$ and $W$ from (20). The simplest case $\theta\equiv 0$ yields the solution
@@
The potential $V^{\rm eaKS}$ is the exact adiabatic Kohn–Sham potential, defined as the total potential for which the time-dependent Kohn–Sham system has ground-state density $\rho^{\rm S}(t)$ for all $t$, and $W^{\text{na}}$ is the corresponding non-adiabatic geometric correction. We emphasize that we here apply the adiabatic approximation to the total exact Kohn–Sham potential $V^{\rm eKS}:=V_{0}[\Phi_{0},\rho^{\rm S}]$. It is not the usual adiabatic approximation, normally applied only to the Hxc potential in (16) [10, Sec. VIII.D]. Nevertheless, it is remarkable that $W^{\text{na}}$ describing all the non-adiabatic effects is such a simple and explicit functional of $\rho^{\rm S}$.
@@
We next consider a two-electron singlet in 1D, with a soft-Coulomb interaction $V_{\text{ee}}(x)=1/\sqrt{x^{2}+1}$. This framework was extensively studied [33, 34, 35, 28, 36, 37], because standard adiabatic TDDFT approximations fail to reproduce, even qualitatively, the exact time-dependent density $\rho^{\rm S}$ for this 1D toy model in much the same way as they do for real 3D Coulombic systems. We consider two different time-dependent external potentials $V_{\rm ext}^{(1)}$ and $V_{\rm ext}^{(2)}$.
First, as in [33, 37], we study Rabi oscillations in a Helium-like atom perturbed by an electric dipole interaction, corresponding to
@@
To find a good approximation to $\rho^{\rm S}$, we follow common practice and solve a two-level time-dependent Schrödinger equation in the basis $(\Psi_{\rm g},\Psi_{\rm e})$. Those are obtained by diagonalizing a finite-difference approximation of the two-electron Hamiltonian at time $t=0$. The non-adiabatic contribution $V^{\text{na}}$ to the time-dependent exact Kohn–Sham potential is computed from (20) with $W=0$. In 1D, we have the explicit formula [38, 39]
@@
Figure 2: Rabi oscillations in a 1D Helium-like atom with soft-Coulomb potential subjected to a time-dependent uniform electric field with amplitude $E=0.00667$ and angular frequency $\omega=0.5336$. Heatmaps of the non-adiabatic contribution $V^{\text{na}}$ to the exact time-dependent Kohn–Sham potential (left) and of the non-adiabatic geometric correction term $W^{\text{na}}$ (right), at times $|t-T_{\rm R}/4|\leqslant 2T_{\rm opt}$ (bottom) and $|t-T_{\rm R}/2|\leqslant 2T_{\rm opt}$ (top) with $T_{\rm opt}=2\pi/\omega$ the optical period.
Figure 3: Charge transfer in 1D between a Helium-like atom and a model of a closed-shell atom with frozen electrons with soft-Coulomb potential obtained by a time-dependent uniform electric field with amplitude $E=0.00667$ and angular frequency $\omega=0.112$. Heatmaps of the non-adiabatic contribution $V^{\text{na}}$ to the exact time-dependent Kohn–Sham potential (left) and of the non-adiabatic geometric correction term $W^{\text{na}}$ (right), at times $|t-T_{\rm R}/8|\leqslant 2T_{\rm opt}$ (bottom) and $|t-T_{\rm R}/4|\leqslant 2T_{\rm opt}$ (top) with $T_{\rm opt}=2\pi/\omega$ the optical period.
@@ Conclusion.
We have proposed a new TDDFT scheme based on geometric considerations. Our numerical simulations for the two-electron singlet suggest that approximating the non-adiabatic correction $W^{\rm na}$ could be easier than the corresponding non-adiabatic part of the standard Kohn–Sham potential, because it is a much smoother function of both time and space. Our new scheme could thus be more adapted to the description of out-of-equilibrium processes. In the companion article [10] we study the mathematical foundations of our geometric approach for finite-dimensional systems. Future work will be devoted to constructing practical approximations of the geometric correction $W$ for real systems.
@@ Acknowledgments.
This work has benefited from French State support managed by ANR under the France 2030 program through the MaQuI CNRS Risky and High-Impact Research programme (RI)2 (grant agreement ANR-24-RRII-0001). We thank the referees for useful comments.
@@ Appendix A Geometric Runge–Gross theorem
The Runge–Gross theorem [1] asserts that two time-dependent external potentials giving rise to the same density must necessarily be equal, up to a time-dependent additive constant. This is a fundamental result in TDDFT. Thus far, the mathematical proof of this result has only been possible under the assumption that the time-dependent potentials are spatially smooth and analytic in time and that the initial wavefunction is smooth [41]. The Runge–Gross argument therefore only applies to very smooth potentials such as the soft-Coulomb potential. Here we provide a rigorous proof that the same arguments as in [41] apply to our geometric theory. More precisely, we work in space dimension $d\in\{1,2,3\}$ and, to cover all situations of interest, we give ourselves four functions $V_{1}(t,\mathbf{r})$, $V_{2}(t,\mathbf{r})$, $W_{1}(t,\mathbf{r})$ and $W_{2}(t,\mathbf{r})$. We will suppose that either $V_{1}=V_{2}$ or $W_{1}=W_{2}$ and want to deduce the missing equality. We will always assume that
  • (i)
@@
  • (ii)
    the functions $V_{1}(t,\mathbf{r})$, $V_{2}(t,\mathbf{r})$, $W_{1}(t,\mathbf{r})$ and $W_{2}(t,\mathbf{r})$ are infinitely differentiable in space-time, with bounded space derivatives for all $t$;
  • (iii)
    the initial state $\Psi_{0}$ is infinitely differentiable in space and all its derivatives are square-integrable.
We also need a further assumption that will be different depending on the situation. We will require that the initial state $\Psi_{0}$ satisfies one of the two conditions:
  • (iv)
    $\rho_{\Psi_{0}}(\mathbf{r})>0$ everywhere;
  • (iv’)
    the unique $V$-representability property
@@
The equalities in (22) are understood Lebesgue-almost everywhere. The assumptions (i)–(iv) are from [41]. The unique $V$-representability property (22) in (iv’) is slightly stronger than (iv) and is further studied in [42] and in [10, Sec. VII.B]. This is the main property used in the proof of the Hohenberg–Kohn theorem for ground-state DFT, where it follows from the unique continuation principle [43, 44]. In particular, (22) is valid if $\Psi_{0}$ is the ground state of a Hamiltonian of the above form. For the theorem we do not need that $\Psi_{0}$ is a ground state but we do assume the condition (22). As explained in [10], this is the main property necessary for the geometric interpretation of TDDFT.
@@
$i\partial_{t}\Psi_{m}(t)=\bigg(\widehat{T}+\widehat{V}_{\rm ee}+\sum_{j=1}^{N}V_{m}(t,\mathbf{r}_{j})+iW_{m}(t,\mathbf{r}_{j})\bigg)\Psi_{m}(t)$
with $m=1$ or $m=2$, starting at the same initial condition $\Psi_{0}$. We assume (i)–(iii) and that $\rho_{\Psi_{1}(t)}=\rho_{\Psi_{2}(t)}$ for $t\in[0,T)$.
  • (a)
    Assume that $W_{1}=W_{2}$, that (iv) holds and that the function $t\mapsto V_{2}(t,\mathbf{r})-V_{1}(t,\mathbf{r})$ is real-analytic for any $\mathbf{r}\in\mathbb{R}^{d}$. Then $V_{1}(t,\mathbf{r})=V_{2}(t,\mathbf{r})+C(t)$ for all $\mathbf{r}\in\mathbb{R}^{d}$ and all $t\in[0,T)$.
  • (b)
    Assume that $V_{1}=V_{2}$, that (iv’) holds and that the function $t\mapsto W_{2}(t,\mathbf{r})-W_{1}(t,\mathbf{r})$ is real-analytic for any $\mathbf{r}\in\mathbb{R}^{d}$. Then $W_{1}(t,\mathbf{r})=W_{2}(t,\mathbf{r})$ for all $\mathbf{r}\in\mathbb{R}^{d}$ and all $t\in[0,T)$.
We implicitly assume in the statement that $\Psi_{m}(t)$ stays normalized for all $t\in[0,T)$, that is, $\int_{\mathbb{R}^{d}}W_{m}(t,\mathbf{r})\rho_{\Psi_{m}(t)}(\mathbf{r})\,\mathrm{d}\mathbf{r}=0$. If we use the commutator form as in (9) then this implicit assumption is not necessary and uniqueness for $W$ holds only up to a time-dependent constant $C(t)$ like for $V$.
The proof of Theorem 1 follows closely [41] and we only explain it in case (b). Under the conditions (i)–(iii) and (iv’), we show that
@@
for all $\ell$ and all $\mathbf{r}\in\mathbb{R}^{d}$. The conclusion then follows from the assumed analyticity in time. The smoothness of the potentials and $\Psi_{0}$ implies that we can freely differentiate Schrödinger’s equation in time [41]. We denote by $\widehat{W}_{m}(t):=\sum_{j=1}^{N}W_{m}(t,\mathbf{r}_{j})$ the many-body operators associated with $W_{1}$ and $W_{2}$. We also introduce the difference $W:=W_{2}-W_{1}$ which we want to prove vanishes, and the corresponding $\widehat{W}:=\widehat{W}_{2}-\widehat{W}_{1}$. We call $\rho(t,\mathbf{r})$ the common density of $\Psi_{1}(t)$ and $\Psi_{2}(t)$. Since in (b) we assume $V_{1}=V_{2}$, we call $H(t)=\widehat{T}+\widehat{V}_{\rm ee}+\widehat{V}_{1}(t)$ the common part of the Hamiltonian.
We then give ourselves a smooth time-independent function $U(\mathbf{r})$, that we will specify later. Using again the notation $\widehat{U}:=\sum_{j=1}^{N}U(\mathbf{r}_{j})$, we have
$\int_{\mathbb{R}^{d}}U(\mathbf{r})\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{U}\Psi_{m}(t)\right\rangle}.$
@@
$\int_{\mathbb{R}^{d}}U(\mathbf{r})\partial_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}\\ =\Big\langle\Psi_{m}(t),\big(i[\widehat{T},\widehat{U}]+\widehat{U}\widehat{W}_{m}(t)+\widehat{W}_{m}(t)\widehat{U}\big)\Psi_{m}(t)\Big\rangle.$ (24)
We evaluate (24) at time $t=0$ and recall that $\Psi_{1}(0)=\Psi_{2}(0)=\Psi_{0}$. Subtracting the two equations (for $m=1$ and $m=2$) leads to
${\left\langle\Psi_{0},(\widehat{U}\widehat{W}(0)+\widehat{W}(0)\widehat{U})\Psi_{0}\right\rangle}=0$
where we recall that $W(0,\mathbf{r})=W_{2}(0,\mathbf{r})-W_{1}(0,\mathbf{r})$. It is natural to take $U(\mathbf{r})=W(0,\mathbf{r})$ so that the previous relation reduces to $\|\widehat{W}(0)\Psi_{0}\|^{2}=0$, that is, $\widehat{W}(0)\Psi_{0}=0$. From the unique $V$-representability property (22), this implies $W(0,\mathbf{r})=W_{2}(0,\mathbf{r})-W_{1}(0,\mathbf{r})=0$ almost everywhere, hence everywhere by continuity. This is the claimed result (23) for $\ell=0$.
Next we go on and differentiate once more (24), still with a general function $U$. To simplify the expression, we denote by $\widehat{L}_{m}^{(1)}(t):=i[\widehat{T},\widehat{U}]+\widehat{U}\widehat{W}_{m}(t)+\widehat{W}_{m}(t)\widehat{U}$ the Hermitian operator appearing in (24). The second-order derivative equals
$\int_{\mathbb{R}^{d}}U(\mathbf{r})\partial^{2}_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{K}^{(1)}_{m}(t)\Psi_{m}(t)\right\rangle}\\ +{\left\langle\Psi_{m}(t),\big(\widehat{U}\partial_{t}\widehat{W}_{m}(t)+\partial_{t}\widehat{W}_{m}(t)\widehat{U}\big)\Psi_{m}(t)\right\rangle}$ (25)
@@
${\left\langle\Psi_{0},\big(\widehat{U}\partial_{t}\widehat{W}(0)+\partial_{t}\widehat{W}(0)\widehat{U}\big)\Psi_{0}\right\rangle}=0.$
Taking $U(\mathbf{r})=\partial_{t}W(0,\mathbf{r})=\partial_{t}W_{2}(0,\mathbf{r})-\partial_{t}W_{1}(0,\mathbf{r})$ and using again (22) for that $U$, we arrive at the conclusion that $\partial_{t}W_{2}(0,\mathbf{r})=\partial_{t}W_{1}(0,\mathbf{r})$, which is (23) for $\ell=1$.
@@
$\int_{\mathbb{R}^{d}}U(\mathbf{r})\partial^{\ell}_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{K}^{(\ell-1)}_{m}(t)\Psi_{m}(t)\right\rangle}\\ +{\left\langle\Psi_{m}(t),\big(\widehat{U}\partial^{\ell}_{t}\widehat{W}_{m}(t)+\partial^{\ell}_{t}\widehat{W}_{m}(t)\widehat{U}\big)\Psi_{m}(t)\right\rangle}$ (26)
@@
$\widehat{L}^{(\ell-1)}_{m}(t):=\widehat{K}^{(\ell-1)}_{m}(t)+\widehat{U}\partial^{\ell-1}_{t}\widehat{W}_{m}(t)+\partial^{\ell-1}_{t}\widehat{W}_{m}(t)\widehat{U}$.